OpenStudy (anonymous):

HELP PLEASE What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

(-2-6)/(1-3)=4=m 1=4(-1)+b 1=-4+b 5=b y=4x+5

OpenStudy (triciaal):

slope = change in y/change in x (-2-6)/(1-3) = -8/-2 =4 parallel lines have the same slope using the point (-1 1) y = mx + b 1 = 4(-1) + b b=5 y = 4x + 5